# A longitudinal Bayesian framework for estimating causal dose-response relationships

### v1

@@ 1 Introduction
The SARS-CoV-2 virus pathogen, which causes the COVID-19 disease, is considered to transmit via two main pathways: (i) directly through droplets and aerosols, and (ii) indirectly through fomite transmission (for example, Derqui et al., 2023). Public transport modes serving dense urban areas induce prolonged exposure for passengers in enclosed and oftentimes poorly ventilated spaces, and they also facilitate the mixing of populations at the origin and destination trip-ends. Consequently, we can reasonably hypothesise that public transport is a vector of virus transmission in a city. While this hypothesis seems intuitively reasonable, and has been put forward previously, including in the mass media (Thomas et al., 2022), it remains largely unsubstantiated statistically under the causal lens. Existing empirical studies quantifying the dynamic impact of mass public transport on COVID-19 transmission are scarce. Most previous studies have analyzed aggregate human mobility, rather than mass transit ridership specifically, for instance, using less granular mobile phone data sources such as those available via Google Mobility or Apple Maps. Despite growing evidence of multi-layered causal pathways, the causal dynamics connecting mass transit ridership, such as metro ridership, to infectious-disease spread remain underexplored.
@@
Figure 1: Longitudinal causal DAG between the mass of metro ridership (treatment variable, $D$), confounders ($X$) and the number of COVID-19 cases ($Y$) from time $t$ to $t+1$. $U$ represents the source of time-invariant unobserved confounding.
[…]
@@
Using the standard causal inference notation, we define our problem in relation to the triple $(Y,D,X)$, indexed by $t$ to denote time: where $Y$ is the outcome (COVID-19 cases in the city), $D$ is the treatment (urban mass transit ridership), $X$ is a set of time-varying confounders, and $U$ is a set of confounders that are time-invariant. We are interested in the causal effects of $D$ on $Y$. Figure 1 shows a directed acyclic graph (DAG) of our assumed time-varying causal relationships between mass transit ridership and disease propagation at the city level. The figure shows an autoregressive structure in $Y$, along with contemporaneous and lagged structure in both treatment effect and confounding. Therefore, we face two key challenges in estimating this causal effect. First, we must adjust for the existence of confounding influences, $X$ and $U$, which determine both transit ridership and disease propagation. These include time-varying (i.e., $X$) factors such as the severity of containment measures on human mobility (e.g. lockdown measures), urban mobility characteristics, and vaccination rates, along with a set of unmeasured time-invariant (i.e., $U$) city level attributes such as characteristics of the built-environment, urban culture, and baseline socio-demographics. Second, there are dynamic effects presented such that past outcomes act as contemporaneous confounders. Thus, we observe a repeatedly measured outcome, a time-dependent treatment and a set of dynamic time-varying and time-invariant confounders. For estimation, we have monthly (longitudinal) city level data available which measure Covid cases, mass transit ridership, mobility containment measures, and a set of mobility characteristics for the period between 1 March 2020 and 23 January 2022. Longitudinal data provide the temporal resolution needed to understand how public transit ridership and COVID-19 incidence evolve and intersect. To determine causality between the repeatedly measured COVID-19 cases and public transportation ridership, we thus require a robust model that explicitly builds the temporal dependencies illustrated in Figure 1.
@@
Most existing methods for causal estimation of a continuous-exposure are developed for the point-treatment framework, including the generalized propensity score-based (GPS) approaches (Robins et al., 2000; Hirano and Imbens, 2004; Galvao and Wang, 2015), g-computation and outcome regression-based approaches (Hill, 2011), and a doubly robust estimator considering combining both treatment and outcome models (Kennedy et al., 2017). Extensions to longitudinal causal data are very limited and restricted to the parametric frequentist paradigms (Robins et al., 2000; Vansteelandt and Sjolander, 2016). The Bayesian lens for longitudinal causal dose-response analysis is largely missing and remains underexplored in the literature. The Bayesian estimation framework can offer both statistical and practical advantages, such as the propagation of uncertainty through complex model structures and the ability to provide probabilistic summaries of causal effects for intuitive causal decision-making in applied research. In addition, it allows for the incorporation of subjective prior beliefs about the treatment effect and confounding from prior scientific knowledge.
@@
In this paper, we fill this literature gap by presenting a flexible yet computationally tractable generalized double non-parametric Bayesian estimation framework that estimates the entire longitudinal dose-response surface while making minimal distribution assumptions about the functional form of the continuous exposure. Our specific methodological contributions in this paper are on three fronts: i) we introduce the generalized double non-parametric Bayesian estimation framework for dose-response causal analysis with longitudinal data, integrating a similar moment-based restrictions to generalized estimating equations with Bayesian uncertainty propagation; ii) we demonstrate the validity of our approach by showing the theoretical basis behind how the approximate frequentist estimation has a natural Bayesian analogue that rests on equally weak assumptions; and iii) we provide computationally scalable algorithms and open-source code making the proposed Bayesian approach practical for large, real-world datasets for wide adoption and application beyond the motivating study.
@@ 2.1 Causal inference for longitudinal data with repeated outcomes
In the causal inference setting, let $Z_{i}=(Y_{i},D_{i},X_{i}),i=1,\ldots,n$, where for the $i$th unit of observation $Y_{i}$ denotes a response, $D_{i}$ the treatment received which can take values in some bounded interval, $\mathcal{D}$ in $\mathbb{R}$, and $X_{i}$ a vector of pre-treatment covariates or confounder variables. Under an experimental design $\mathcal{E}$, we know that $D_{i}\perp\!\!\!\perp_{\mathcal{E}}X_{i}$ with the marginal distribution of $D_{i}$, $f_{\mathcal{E}}(d_{i})$ known as a constant. However, in observational studies, $\mathcal{O}$, $D_{i}\not\perp\!\!\!\perp_{\mathcal{O}}X_{i}$ and an appropriate adjustment is required to estimate causal effects. Under the potential outcome framework, the observed outcome is, $Y_{i}=\sum_{d\in\mathcal{D}}\mathbbm{1}_{(D_{i}=d)}Y_{i}(d)$, where $\mathcal{Y}_{i}=\left\{Y_{i}(d):d\in\mathcal{D}\right\}$, is set of all potential outcomes. We assume that the treatment assignment is ignorable, $D_{i}\perp\!\!\!\perp_{\mathcal{O}}\mathcal{Y}_{i}\left|X_{i}\right.$, and we also assume that the potential outcomes $\mathcal{Y}_{i}$ are not affected by any other subjects, which is termed the stable unit-treatment value assumption (SUTVA). The primary focus of our inference is average potential outcomes (APOs) for a population under $\mathcal{O}$, i.e., $\mathbb{E}\left[Y_{i}(d)\right]$, given a specific dose/treatment $d$. In the longitudinal setting, we wish to study the APO of a longitudinal exposure pattern, over $K$ intervals, on a repeatedly measured outcome. We have for $i=1,\ldots,n$, the observed data can be written as $Z_{i}=\left(X_{i1},D_{i1},Y_{i1},X_{i2},D_{i2},Y_{i2},\ldots,X_{iK},D_{iK},Y_{iK}\right)$, where $D_{ik}$ is the treatment variable for individual $i$ at time $k$, $X_{ik}$ is the confounder measured prior to the treatment at time $k$ for individual $i$, and $Y_{ik}$ is the corresponding outcome.
We denote the observed covariate history $\bar{x}_{k}$ as $\bar{x}_{k}=\left(x_{1},x_{2},\ldots,x_{k}\right)$. Following Robins et al. (2000), we assume that $Y(d_{k})\perp\!\!\!\perp D_{k}|\bar{X}_{k}$, which is equivalent to the assumption of no unmeasured confounders in the cross-section analysis. Based on the observable, the joint distribution for individual $i$ is
$f\left(z_{i}|\xi\right)=\prod_{k=1}^{K}f_{\mathcal{O}}\left(y_{ik}\left|d_{ik},x_{ik};\xi\right.\right)f_{\mathcal{O}}\left(d_{ik}\left|x_{ik};\gamma\right.\right)f_{\mathcal{O}}\left(x_{ik}\right)$ (1)
@@
where $f_{\mathcal{O}}\left(d_{ik}\left|x_{ik};\gamma\right.\right)\equiv e(x_{ik};\gamma)$ represents the PS model and $f_{\mathcal{O}}\left(y_{ik}\left|d_{ik},x_{ik};\xi\right.\right)$ is the outcome mean model. In line with standard practice in the causal inference, we do not aim to model the marginal distribution $f_{\mathcal{O}}\left(x_{ik}\right)$. The PS in this case will retain the balancing property, that is, $X_{k}\perp\!\!\!\perp D_{k}|e(x_{k};\gamma)$. Graham et al. (2014) proposed a longitudinal mixed model extension of the generalized propensity score estimator to handle time-varying confounding and bi-directionality between treatment and outcome. The two-step approach first uses a random intercept in the marginal PS model to adjust for time-invariant confounding, followed by an outcome regression model incorporating the estimated PS. This procedure yields consistent estimates of the APO if the PS models are correctly specified. Building on this approach, we can model the marginal outcome distribution by incorporating the estimated PS. The marginal structural model uses subject-visit-specific weights to create a pseudo-population in which treatment is independent of covariates and outcome history (Liu et al., 2020). The marginal treatment effect can then be estimated using a generalized estimating equation (GEE), defined by the following estimating equation
@@
$\sum_{i=1}^{n}\mathbf{U}_{i}\left(\xi\right)={\sum\limits_{i=1}^{n}{\left({{{\partial{u}(x_{i},d_{i},e(x_{i},\widehat{\gamma}),\xi)}\over{\partial\xi}}}\right)}^{\top}}\Sigma_{i}^{-1}\left({{y_{i}}-{u}(x_{i},d_{i},e(x_{i},\widehat{\gamma}),\xi)}\right)=0$ (2)
@@
where ${{y_{i}}}$ is a vector of observed response for $ith$ subject. The function ${u_{i}}(x_{i},d_{i},e(x_{i},\widehat{\gamma}),\xi)$ is the marginal conditional mean model of ${{y_{i}}}$ with the estimate PS via a longitudinal model, such as a mixed effect model (Graham et al., 2014) or a GEE (Liu et al., 2020). This connects the targeting parameter $\xi$ from a link function, and $\Sigma_{i}$ is the $(K+1)\times(K+1)$ working covariance matrix for subject $i$. However, as noted in Tchetgen et al. (2012), the GEE approach consistently estimates the parameters of a correctly specified regression model in absence of weights, irrespective of whether the working correlation structure is correct. The GEE framework offers flexibility in targeting the marginal treatment effect for repeated measurements. However, its reliance on estimating equations, rather than full likelihoods, poses challenges for conventional Bayesian analysis, as there is no explicit distributional form for the marginal mean model. This motivates the development of a fully Bayesian approach, which we outline in the next section.
@@ 2.2 Bayesian non-parametric method
[…]
@@ 2.3 The generalized Bayesian bootstrap
$\sum_{j=1}^{n}p_{j}\mathbf{U}_{1}\left(x_{j},d_{j},\gamma\right)$ $=\mathbf{0}$ (5)
$\sum_{j=1}^{n}p_{j}\mathbf{U}_{2}\left(z_{j},\xi,e(x_{j},\widehat{\gamma})\right)$ $=\mathbf{0},$
where the GEE, $\mathbf{U}_{1}(\cdot)$, is the estimating equation related to the propensity score model, and $\mathbf{U}_{2}(\cdot)$ is the GEE for the OR model, and $\widehat{\gamma}$ is the solution to (5). Therefore, we can obtain the posterior distribution of $\xi$ under this limiting specification (Newton and Raftery, 1994; Chamberlain and Imbens, 2003).
@@
$\sum_{j=1}^{\infty}p_{j}\mathbf{U}_{1}\left(x^{s}_{j},d^{s}_{j},\gamma\right)=\mathbf{0},\sum_{j=1}^{\infty}p_{j}\mathbf{U}_{2}\left(z^{s}_{j},\xi,e(x_{j},\widehat{\gamma})\right)=\mathbf{0}.$ (6)
Although this is an infinite sum, the $p_{j}$ decreases in expectation as $j$ increases and eventually becomes numerically negligible. Because of this decay, the infinite sum can be truncated at some finite $J$, such that $p_{J}<\epsilon$ for a small tolerance $\epsilon$. The uncertainty about $\xi$ is properly addressed in this ‘forward’ calculation. We also layout this DP specification for the longditual analysis in Algorithm 1. In this formulation, we do not directly specify a prior for $\xi$ but instead it is a functional of $p_{i}$, where we carry out a fully Bayesian inference for $p_{i}$. This approach is computationally more efficient because it relies purely on optimization instead of MCMC, which provides a ‘shortcut’ to fully Bayesian causal inference. The validity and theoretical justification of this approach have been thoroughly examined in in cross-sectional settings in Luo et al. (2023).
@@
  • Sample a new collection of data$\{z^{s}_{k}\}$ with probability $1/(\alpha+n)$ from $z_{i}$ and probability $\alpha/(\alpha+n)$ from $G_{0}$;
@@
  • Estimate the PS based the GEE related to $\mathbf{U}_{1}$ in (6) with the data $\{z^{s}_{k}\}$ and the weight $\{p_{k}^{s}\}$;
  • Solve the GEE related to $\mathbf{U}_{2}$ in (6) and obtain $\xi^{s}$ with the data $\{z^{s}_{k}\}$, the weight $\{p_{k}^{s}\}$ and the estimated PS;

### v2

@@ 1 Introduction
The SARS-CoV-2 virus pathogen, which causes the COVID-19 disease, is considered to transmit via two main pathways: (i) directly through droplets and aerosols, and (ii) indirectly through fomite transmission (for example, Derqui et al., 2023). Public transport modes serving dense urban areas induce prolonged exposure for passengers in enclosed and oftentimes poorly ventilated spaces, and they also facilitate the mixing of populations at the origin and destination trip-ends. Consequently, we can reasonably hypothesize that public transport is a vector of virus transmission in a city. While this hypothesis seems intuitively reasonable, and has been put forward previously, including in the mass media (Thomas et al., 2022), it remains largely unsubstantiated statistically under the causal lens. Existing empirical studies quantifying the dynamic impact of mass public transport on COVID-19 transmission are scarce. Most previous studies have analyzed aggregate human mobility, rather than mass transit ridership specifically, for instance, using less granular mobile phone data sources such as those available via Google Mobility or Apple Maps. Despite growing evidence of multi-layered causal pathways, the causal dynamics connecting mass transit ridership, such as metro ridership, to infectious-disease spread remain underexplored.
@@
Figure 1: Longitudinal causal DAG between the mass of metro ridership (treatment variable, $D$), confounders ($X$) and the number of COVID-19 cases ($Y$) from time $t$ to $t+1$.
$Y_{t}$$D_{t+1}$$X_{t+1}$$X_{t}$$Y_{t+1}$$D_{t}$}
@@
Using the standard causal inference notation, we define our problem in relation to the triple $(Y,D,X)$, indexed by $t$ to denote time: where $Y$ is the outcome (COVID-19 cases in the city), $D$ is the treatment (urban mass transit ridership), $X$ is a set of time-varying confounders We are interested in the causal effects of $D$ on $Y$. Figure 1 shows a directed acyclic graph (DAG) of our assumed time-varying causal relationships between mass transit ridership and disease propagation at the city level. The figure shows an autoregressive structure in $Y$, along with contemporaneous and lagged structure in both treatment effect and confounding. Therefore, we face two key challenges in estimating this causal effect. First, we must adjust for the existence of confounding influences, $X$ and $U$, which determine both transit ridership and disease propagation. These include time-varying (i.e., $X$) factors such as the severity of containment measures on human mobility (e.g. lockdown measures), urban mobility characteristics, and vaccination rates. Second, there are dynamic effects presented such that past outcomes act as contemporaneous confounders. Thus, we observe a repeatedly measured outcome, a time-dependent treatment and a set of dynamic time-varying and time-invariant confounders. For estimation, we have monthly (longitudinal) city level data available which measure Covid cases, mass transit ridership, mobility containment measures, and a set of mobility characteristics for the period between 1 March 2020 and 23 January 2022. Longitudinal data provide the temporal resolution needed to understand how public transit ridership and COVID-19 incidence evolve and intersect. To determine causality between the repeatedly measured COVID-19 cases and public transportation ridership, we thus require a robust model that explicitly builds the temporal dependencies illustrated in Figure 1.
@@
Most existing methods for causal estimation of a continuous-exposure are developed for the point-treatment framework, including the generalized propensity score-based (GPS) approaches (Robins et al., 2000; Hirano and Imbens, 2004; Galvao and Wang, 2015), g-computation and outcome regression-based approaches (Hill, 2011), and a doubly robust estimator considering combining both treatment and outcome models (Kennedy et al., 2017). Extensions to longitudinal causal data are very limited and restricted to the parametric frequentist paradigms (Robins et al., 2000; Vansteelandt and Sjolander, 2016). In particular, g-computation can be computationally intensive and intractable for dose–response estimation in the longitudinal setting with time-varying treatment and confounding and repeatedly measured outcomes, as it requires fitting sequential models of the complete data-generating process and simulating counterfactual trajectories of time-varying covariates and outcomes at a set of treatment levels. (Robins, 1986; Linero, 2022). Therefore, in this paper, we fill this literature gap by presenting a flexible yet computationally tractable generalized double non-parametric Bayesian estimation framework that estimates the entire longitudinal dose-response surface while making minimal distribution assumptions about the functional form of the continuous exposure. The Bayesian lens for longitudinal causal dose-response analysis is largely missing and remains underexplored in the literature. The Bayesian estimation framework can offer both statistical and practical advantages, such as the propagation of uncertainty through complex model structures and the ability to provide probabilistic summaries of causal effects for intuitive causal decision-making in applied research. In addition, it allows for the incorporation of subjective prior beliefs about the treatment effect and confounding from prior scientific knowledge.
@@
Our specific methodological contributions in this paper are on three fronts: i) we introduce the generalized double non-parametric Bayesian estimation framework for dose-response causal analysis with longitudinal data, integrating a similar moment-based restrictions to generalized estimating equations with Bayesian uncertainty propagation; ii) we demonstrate the validity of our approach by showing the theoretical basis behind how the approximate frequentist estimation has a natural Bayesian analogue that rests on equally weak assumptions; and iii) we provide computationally scalable algorithms and open-source code making the proposed Bayesian approach practical for large, real-world datasets for wide adoption and application beyond the motivating study.
@@ 2.1 Causal inference for longitudinal data with repeated outcomes
In causal inference for the longitudinal setting, for $t=0,\ldots,T$, let $D_{t}\in\mathbb{R}$ denote the (possibly continuous) dose at time $t$, $\bar{D}_{t}=(D_{t-c},\ldots,D_{t})$ the treatment dosing history where $1\leq c\leq t-1$ specifies the length of the relevant treatment history. Therefore, $\bar{D}_{t}$ represents the sequence of recent doses up to and including time $t$, which may influence the potential outcome $Y_{t}$. In our COVID analysis, we set $c=1$, corresponding to a first-order dosing history. However, the framework can be generalized to larger values of $c$, allowing the outcome $Y_{t}$ to depend on a longer history of past doses. $X_{t}$ a set of time-varying covariates, and $Y_{t}$ the continuous outcome measured at $t$. Using the potential outcome framework in the longitudinal setting, we set $Y_{t}(\bar{d}_{t})$ to denote the counterfactual outcome at time $t$ that would be observed if an individual were assigned the dose history $\bar{d}_{t}=(d_{1},\ldots,d_{t})$ up to time $t$. The primary focus of our inference is on average potential outcomes (APOs) for a population under a hypothetical intervention, given the dose history $\bar{d}_{T}=(d_{1},\ldots,d_{T})$ over $T$ intervals. That is, the estimand of interest is $\mathbb{E}\!\left[Y_{t}(\bar{d}_{t})\right]$. To identify the APO, we require several assumptions. First, we assume that there is no interference across units, i.e., for each individual $i$, $Y_{t,i}(\bar{d}_{1},\ldots,\bar{d}_{n})=Y_{t,i}(\bar{d}_{i})$. This is termed as the stable unit treatment value assumption (SUTVA). Second, we require that if an individual actually follows dose history $\bar{D}_{t}=\bar{d}_{t}$, then their observed outcome at each time matches the corresponding counterfactual, i.e. $Y_{t}=Y_{t}(\bar{d}_{t})$ for all $t=1,\ldots,T$, so that the observed trajectory equals the counterfactual trajectory under the realized dose sequence. Third, we assume that, at each time $t=1,\ldots,T$, conditional on the covariate $X_{t}$, past doses $\bar{D}_{t-1}$, and past outcomes $\bar{Y}_{t-1}$, the dose assignment at $t$ is independent of the future counterfactual outcomes, i.e. $Y_{t}(\bar{d}_{t})\perp\!\!\!\perp D_{t}\mid X_{t},\bar{D}_{t-1}$. Finally, for any history $\{x_{t},\bar{d}_{t-1}\}$ with positive density, we assume that the conditional density of treatment assignment satisfies $f(d_{t}\mid X_{t}=x_{t})>0$ over the range of doses considered for the target dose–response. Following the standard marginal effect parametrization of time-varying treatment and a repeatedly measured outcome, the APOs at each time $t$ is posited as
$g(\mathbb{E}\!\left[Y_{t}(\bar{d}_{t})\right])=m(\bar{d}_{t},x_{t};\,\xi),$
@@
where $m(\cdot;\xi)$ is the conditional mean function of the treatment history $\bar{d}_{t}$ and the time-varying covariates $x_{t}$ characterized by the set of parameters $\xi$, and $g(\cdot)$ is the link function. Marginal structural models (MSMs) can be applied to directly model the marginal distribution of the outcome, while inverse probability weights are applied at the subject–visit level to construct a pseudo-population in which treatment assignment is independent of past covariates and outcome history. This reweighting removes time-dependent confounding, enabling consistent estimation of causal effects. Instead of reweighting, Graham et al. (2014) proposed a longitudinal mixed model extension of the generalized propensity score estimator to handle time-varying confounding and bi-directionality between treatment and outcome. This approach requires incorporating the generalized propensity score directly into the outcome model, $m(\bar{d}_{t},x_{t};\,\xi)$. This procedure yields consistent estimates of the APO if the propensity score model is correctly specified. Therefore, one can rewrite the MSM with functions of the estimated treatment dose densities to adjust for time-dependent confounding, i.e.,
$g(\mathbb{E}[Y_{t}\mid\bar{D}_{t},X_{t},\widehat{e}_{t}(X_{t})])=m(\bar{D}_{t},X_{t};\xi)+h\!\left(\widehat{e}_{t}(X_{t})\right),$
where $\widehat{e}_{t}(\cdot)$ denotes the estimated generalized propensity score at time $t$, and $h(\cdot)$ is a pre-specified adjustment function, which can be estimated some non-parametric method, such as splines. Therefore, the marginal treatment effect, $\xi$, can then be estimated using a generalized estimating equation (GEE) (Preisser et al., 2002), defined by the following estimating equation
@@
$\sum_{i=1}^{n}\mathbf{U}_{i}\left(\xi\right)={\sum\limits_{i=1}^{n}{\left({{{\partial g(\mathbb{E}[Y_{i}\mid\bar{D}_{i},X_{i},\widehat{e}_{i}(X_{i})])}\over{\partial\xi}}}\right)}^{\top}}\Sigma_{i}^{-1}\left({{Y_{i}}-g(\mathbb{E}[Y_{i}\mid\bar{D}_{i},X_{i},\widehat{e}_{i}(X_{i})])}\right)=0$ (1)
@@
where $Y_{i}$}} is a vector of observed response for ithith subject. This connects the targeting parameter $\xi$ from a link function, and $\Sigma_{i}$ is the $T\times T$ working covariance matrix for subject $i$. However, as noted in Tchetgen et al. (2012), the GEE approach consistently estimates the parameters of a correctly specified outcome model in absence of weights, irrespective of whether the working correlation structure is correct. The GEE framework offers flexibility in targeting the marginal treatment effect for repeated measurements. However, its reliance on estimating equations, rather than full likelihoods, poses challenges for conventional Bayesian analysis, as there is no explicit distributional form for the marginal mean model. This motivates the development of a fully Bayesian approach, which we outline in the next section.
@@ 2.2 Bayesian non-parametric method
The Bayesian framework offers calibrated uncertainty quantification and prediction, which can render a more thorough understanding of the causal effect from a practical perspective. There is growing interest in the application of Bayesian methodology to causal problems, but most proposed methods appeal to the principles of PS-adjustment or flexible outcome modeling (Stephens et al., 2023; Li et al., 2023). An advantage of the Bayesian approach for estimation of these models is that estimated causal effects are in the form of distributions, rather than point estimates, allowing us to make probability statements about the causal quantities of interest (i.e., APOs). For example, we can specify a parametric outcome model and propensity score model and then draw inference based on the posterior sample. This is challenging in the causal setting as the GEE model (1) is only partially specified via the conditional mean without specification of the distribution model. We have to examine procedures for Bayesian inference in the case where we wish to perform analysis of an approximate model, acknowledged to be misspecified compared to the data generating model. Suppose we assume that there is a set of estimating equations $\mathbf{U}\left(\xi\right)$ of functions such that $\mathbb{E}[\mathbf{U}\left(\xi\right)|\xi]=0$, for all $\xi$. The objective is to find a non-parametric approximate model, $p(z)$, to the true data generating model $f(z|\xi)$. Thus, we can define some discrepancy $\delta(f|p)$, subject to $\int\mathbf{U}\left(\xi\right)p(z)dz=0$ and $\int p(z)dz=1$. The problem then becomes
$\min_{p}\delta(f|p)\text{ subject to }\int\mathbf{U}\left(\xi\right)p(z)dz=0,\;\;\int p(z)dz=1,\;\;\forall\xi\in\Xi.$ (2)
@@ 2.3 The generalized Bayesian bootstrap
$\sum_{j=1}^{n}p_{j}\mathbf{U}^{(1)}_{j}\left(x_{j},d_{j},\gamma\right)$ $=\mathbf{0}$ (4)
[…]
where the GEE, $\mathbf{U}^{(1)}(\cdot)$, is the estimating equation related to the propensity score model, and $\mathbf{U}^{(2)}(\cdot)$ is the GEE for the OR model, and $\widehat{\gamma}$ is the solution to (4). Therefore, we can obtain the posterior distribution of $\xi$ under this limiting specification (Newton and Raftery, 1994; Chamberlain and Imbens, 2003).
@@
$\sum_{j=1}^{\infty}p_{j}\mathbf{U}^{(1)}_{j}\left(x^{s}_{j},d^{s}_{j},\gamma\right)=\mathbf{0},\sum_{j=1}^{\infty}p_{j}\mathbf{U}^{(2)}_{j}\left(z^{s}_{j},\xi,e(x_{j},\widehat{\gamma})\right)=\mathbf{0}.$ (5)
Although this is an infinite sum, the $p_{j}$ decreases in expectation as $j$ increases and eventually becomes numerically negligible. Because of this decay, the infinite sum can be truncated at some finite $J$, such that $p_{J}<\epsilon$ for a small tolerance $\epsilon$. The uncertainty about $\xi$ is properly addressed in this ‘forward’ calculation. We also layout this DP specification for the longitudinal analysis in Algorithm 1. In this formulation, we do not directly specify a prior for $\xi$ but instead it is a functional of $p_{i}$, where we carry out a fully Bayesian inference for $p_{i}$. This approach is computationally more efficient because it relies purely on optimization instead of MCMC, which provides a ‘shortcut’ to fully Bayesian causal inference. The validity and theoretical justification of this approach have been thoroughly examined in cross-sectional settings in Luo et al. (2023).
@@
[…]
@@
  • Estimate the PS based the GEE related to $\mathbf{U}^{(1)}$ in (5) with the data $\{z^{s}_{k}\}$ and the weight $\{p_{k}^{s}\}$;
  • Solve the GEE related to $\mathbf{U}^{(2)}$ in (5) and obtain $\xi^{s}$ with the data $\{z^{s}_{k}\}$, the weight $\{p_{k}^{s}\}$ and the estimated PS;